OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80 A. an = -5 • (-4)n B. an = -5(-4)n-1 C. an = -4(-5)n-1 D. an = -5 • (4)n

OpenStudy (anonymous):

Here is another one, I have no idea what this stuff is. Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32 A. an = -4 + an-1 B. an = -2 + an-1 C. an = -2 • an-1 D. an = -4 • an-1 Its the last one on the test, though. So thats good.

OpenStudy (anonymous):

Do you need to know the difference between recursive and explicit formulas?

OpenStudy (anonymous):

Um.... I'm not sure. I'm trying to do this test without the textbook - my school won't let me have it until my teacher grades my final exam on Algebra 1 part 1.

OpenStudy (anonymous):

I don't really care about learning it right now, so long as I can get the correct answer on these. I'll just learn it when I get the textbook.

OpenStudy (anonymous):

ok. To be sure, an explicit formula expresses every term of a sequence or series as some operation on a variable. A recursive formula "returns" to a previous term to find the next one. A good policy is to test each of the given answers by substituting. It can take time, but it's worth it.